Weather Report for October 20, 2015

Weather problems continue at the Kuiper Telescope. The sky cleared for almost three hours at the start of the night. Unfortunately, much of this time was spent working out issues having to do with the new telescope control system that was recently installed. Some standard stars were observed, but by the time the first blazar could be observed, clouds had moved back in and only a very low quality observation could be made. The dome was closed as the humidity rose and the fog moved in. The observatory was in fog for the rest of the night.